OpenStudy (anonymous):

Solve for X: 1 over 3 (2x - 8) = 4 2 6 10 10 over 3

OpenStudy (anonymous):

what do u think it is

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

yass

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

since 4 is a third of the total i multiplied 4 times 3 and got 12 then added 8 and got 20 then divided it by 2

OpenStudy (anonymous):

and got 10

OpenStudy (anonymous):

ok look at this to help you if you don't understand it http://openstudy.com/study#/updates/50e348cbe4b0e36e35142b74

OpenStudy (anonymous):

that one was fairly easier

OpenStudy (anonymous):

i guess i just needed to try harder

OpenStudy (anonymous):

yea good job tho

OpenStudy (anonymous):

thanx i got a 100 on my test and i only asked 2 question!! YAY!

OpenStudy (anonymous):

good lol give me a medal

OpenStudy (texaschic101):

1/3(2x - 8) = 4 --- multiply both sides by 3 to get rid of the fraction 2x - 8 = 12 2x = 12 + 8 2x = 20 x = 10 yep...thats it